. In each part, sketch the graph of the equation in 3-space. (a) x = y 2 (b) z = x 2 (c) y = z 2 


1
Expert's answer
2021-08-09T13:15:23-0400

Given

(a) "x=y^2"

About z-axis




(b) "z=x^2"

About y axis




(c) "y=z^2"

About x-axis